OpenStudy (anonymous):

what is the equation in slope intercept form of the line that has a slope of -3 and contains the point (1,1)

OpenStudy (anonymous):

use the formula (y-y1)=m(x-x1) where you plug in the known ordered pair into(x1,y1)

OpenStudy (anonymous):

i got y-1=-3x-3 but idk what else to do

OpenStudy (anonymous):

when you distribute..it would be a -3*-1 so what should the last number be?

OpenStudy (anonymous):

3

OpenStudy (anonymous):

yupp so instead is would be y-1=-3x+3 then you would just add 1 to each side to get y by itself.